OpenStudy (anonymous):

find the slope for r=6/theta, theta=pi/3

OpenStudy (anonymous):

\[\large\frac{dy}{dx}=\frac{\frac{dy}{d\theta}}{\frac{dx}{d\theta}}\] where \(\dfrac{dy}{dx}\) denotes the slope of the tangent line. Recall that when converting from rectangular to polar coordinates, you have the relations \[\begin{cases}x=r\cos\theta\\y=r\sin\theta\end{cases}\] Keep in mind that \(r\) is a function of \(\theta\). Denote this by \(r=f(\theta)\). \[\begin{cases}x=f(\theta)\cos\theta\\y=f(\theta)\sin\theta\end{cases}\] Then \(f'(\theta)=\dfrac{dr}{d\theta}\). Differentiate both equations with respect to \(\theta\): \[\begin{cases} \dfrac{dx}{d\theta}=f'(\theta)\cos\theta-f(\theta)\sin\theta=\dfrac{dr}{d\theta}\cos\theta-r\sin\theta\\\\ \dfrac{dy}{d\theta}=f'(\theta)\sin\theta+f(\theta)\cos\theta=\dfrac{dr}{d\theta}\sin\theta+r\cos\theta \end{cases}\] which gives \[\large\frac{dy}{dx}=\frac{\frac{dr}{d\theta}\cos\theta-r\sin\theta}{\frac{dr}{d\theta}\sin\theta+r\cos\theta}\]